OpenStudy (anonymous):

If A= [6 -1], & B= [-4 2], Find A-B [3 -8] [-2 3] [3 -4] [0 1]

OpenStudy (callisto):

A-B \[=\left[\begin{matrix}6 & -1 \\ 3 & -8 \\ 3 & -4\end{matrix}\right] - \left[\begin{matrix}-4 & 2 \\ -2 & 3 \\ 0 & 1\end{matrix}\right]\]\[=\left[\begin{matrix}6-(-4) & -1-2 \\ 3-(-2) & -8-3 \\ 3-0 & -4-1\end{matrix}\right]\]\[=\left[\begin{matrix}6+4 & -1-2 \\ 3+2 & -8-3 \\ 3 & -4-1\end{matrix}\right]\]\[=...?\]

OpenStudy (anonymous):

@cheer_bby18 Knock, knock are you there?